# Many-body Fu-Kane-Mele index

### v1

@@ B.1 Locality estimates for Fermi projections
The trace class claim follows immediately. ∎
B.2 The spin Hall Chern number as an index of a pair of projections
By Proposition B.5 the difference ${\mathscr{P}}_{+}-{\mathscr{P}}_{-}$ is trace class so the projections ${\mathscr{P}}_{+}$ and ${\mathscr{P}}_{-}$ have a well-defined index of a pair of projections which is given by
$\mathrm{ind}({\mathscr{P}}_{+},{\mathscr{P}}_{-})=\operatorname{Tr}\{{\mathscr{P}}_{+}-{\mathscr{P}}_{-}\}.$ (B.32)
We will show that this index is equal to the Chern number of the Fermi projection ${\mathscr{P}}$. This Chern number may be expressed as (see for example Eq. 1.4 of [19])
$\mathrm{Ch}({\mathscr{P}})=2\pi{\mathrm{i}}\operatorname{Tr}\{{\mathscr{P}}[[\Pi_{h},{\mathscr{P}}],[\Pi_{R},{\mathscr{P}}]]\}$ (B.33)
[…]
Let $\Pi(r)$ be the projection onto the box $\{(x_{1},x_{2})\in{\mathbb{Z}}^{2}\,:\,\left|x_{1}\right|,\left|x_{2}\right|\leq r\}$, then
$\mathrm{ind}({\mathscr{P}}_{+},{\mathscr{P}}_{-})$ $=\lim_{r\to\infty}\operatorname{Tr}\{\Pi(r)({\mathscr{P}}_{+}-{\mathscr{P}}_{-})\Pi(r)\}$ (B.34)
Since both terms are now separately trace class,
$=-{\mathrm{i}}\lim_{r\to\infty}\int_{-\pi}^{\pi}{\mathrm{d}}\phi\operatorname{Tr}\{{\mathscr{P}}_{\phi}[{\mathscr{K}}_{\phi},\Pi(r)]\}$ (B.35)
using repeatedly the cyclicity of the trace. Since ${\mathscr{K}}_{\phi}$ is localised near the left horizontal axis, where it is equal to ${\mathscr{K}}_{\phi}^{h}$ (Lemma B.1 and ${\mathscr{K}}_{\phi}=\Pi_{L}{\mathscr{K}}^{h}_{\phi}\Pi_{L}$) this becomes
$=-{\mathrm{i}}\lim_{r\to\infty}\int_{-\pi}^{\pi}{\mathrm{d}}\phi\operatorname{Tr}\{{\mathscr{P}}_{\phi}[{\mathscr{K}}^{h}_{\phi},\Pi_{\geq-r}]\}$ (B.36)
where $\Pi_{\geq-r}$ is the projection onto the half-plane $\{(x_{1},x_{2})\in{\mathbb{Z}}^{2}\,:\,x_{1}\geq-r\}$. Similarly since ${\mathscr{P}}_{\phi}-{\mathscr{P}}^{h}_{\phi}$ is localised on the right half axis (immediate consequence of the first bound of Lemma B.4) and ${\mathscr{K}}_{\phi}^{h}=-{\mathrm{i}}[\partial_{\phi}{\mathscr{P}}_{\phi}^{h},{\mathscr{P}}_{\phi}^{h}]$,
$=-\lim_{r\to\infty}\int_{-\pi}^{\pi}{\mathrm{d}}\phi\operatorname{Tr}\{{\mathscr{P}}^{h}_{\phi}[[\partial_{\phi}{\mathscr{P}}^{h}_{\phi},{\mathscr{P}}^{h}_{\phi}],\Pi_{\geq-r}]\}.$ (B.37)
Jacobi’s identity reads
$-[[\partial_{\phi}{\mathscr{P}}^{h}_{\phi},{\mathscr{P}}^{h}_{\phi}],\Pi_{\geq-r}]=[[{\mathscr{P}}^{h}_{\phi},\Pi_{\geq-r}],\partial_{\phi}{\mathscr{P}}^{h}_{\phi}]+[[\Pi_{\geq-r},\partial_{\phi}{\mathscr{P}}^{h}_{\phi}],{\mathscr{P}}^{h}_{\phi}].$ (B.38)
Since ${\mathscr{P}}^{h}_{\phi}$ is translation invariant in the horizontal direction, the first term is a translate of $[[{\mathscr{P}}^{h}_{\phi},\Pi_{R}],\partial_{\phi}{\mathscr{P}}^{h}_{\phi}]$ and can therefore be replaced by the latter in the trace. The second term is off-diagonal with respect to ${\mathscr{P}}^{h}_{\phi}$ so its contribution vanishes in the trace. Noting that the derivative is given by $\partial_{\phi}{\mathscr{P}}^{h}_{\phi}={\mathrm{i}}[\Pi_{h},{\mathscr{P}}^{h}_{\phi}]$ we therefore obtain
$\mathrm{ind}({\mathscr{P}}_{+},{\mathscr{P}}_{-})={\mathrm{i}}\int_{-\pi}^{\pi}{\mathrm{d}}\phi\operatorname{Tr}\{{\mathscr{P}}^{h}_{\phi}[[{\mathscr{P}}^{h}_{\phi},\Pi_{R}],[\Pi_{h},{\mathscr{P}}^{h}_{\phi}]]\},$ (B.39)
and since the integrand does not depend on $\phi$ (by unitary invariance of the trace and $[\Pi_{h},\Pi_{R}]=0$), we conclude that
$\mathrm{ind}({\mathscr{P}}_{+},{\mathscr{P}}_{-})=-2\pi{\mathrm{i}}\operatorname{Tr}\left\{{\mathscr{P}}\left[[\Pi_{R},{\mathscr{P}}],[\Pi_{h},{\mathscr{P}}]\right]\right\}.$ (B.40)
Recalling that we wrote ${\mathscr{P}}={\mathscr{P}}^{\uparrow}$ and ${\mathscr{P}}_{\pm}={\mathscr{P}}_{\pm}^{\uparrow}$, and comparing with  (B.33) we have proven
Proposition B.6.
The difference ${\mathscr{P}}_{+}^{\uparrow}-{\mathscr{P}}_{-}^{\uparrow}$ is trace class and the Chern number $\mathrm{Ch}({\mathscr{P}}^{\uparrow})$ of ${\mathscr{P}}^{\uparrow}$ is equal to the index of a pair of projections $\mathrm{ind}({\mathscr{P}}^{\uparrow}_{+},{\mathscr{P}}^{\uparrow}_{-})$.

### v2

@@ C.3 Comparison of quasi-adiabatic and spectral evolutions
The fact that this difference is trace class follows immediately from
${\mathscr{P}}_{\phi}^{(\mu)}={\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})}{\mathscr{P}}_{\phi^{(i)}}^{(\mu)}({\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})})^{*}$
and Lemma C.5. The trace of the difference is the index of the pair of projections, which is also the Fredholm index of the Fredholm operator ${\mathscr{P}}_{\phi^{(i)}}^{(\mu)}({\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})})^{*}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}{\mathscr{P}}_{\phi^{(i)}}^{(\mu)}$ seen as an operator on $\mathrm{Ran}{\mathscr{P}}_{\phi^{(i)}}^{(\mu)}$ (Theorem 5.2 of [7]). Since the index of a Fredholm operator which is a compact perturbation of the identity vanishes, the vanishing of the trace follows again from Lemma C.5. ∎